If a 1.25 g sample zinc oxide, 95% ZnO, were treated with 50mL of 1.1230N sulfuric acid

in the usual way, what volume of 0.9765N sodium hydroxide would be required in the

back titration?


1
Expert's answer
2021-03-24T01:06:08-0400

ZnO + H2SO4 = ZnSO4 + H2O

n(H2SO4) = n(ZnO) = 0.95*(1.25 g)/(81.38 g/mol) = 0.0146 mol

2NaOH + H2SO4 = Na2SO4 + 2H2O

Molarity of H2SO4 if half its normality, so

n(NaOH) = 2*n(H2SO4)left = 2*((0.05 L)*(1.1230/2 M) - 0.0146 mol) = 0.02695 mol

Molarity of NaOH if the same as its normality, so

V(NaOH) = (0.02695 mol)/(0.9765 M) = 0.0276 L = 27.6 mL